# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v1

@@ 3 The Proposed Metrics
CI-Dep: We propose a modified version of Harrell’s CI [Harrell Jr et al., 1982] that employs inverse probability weighting (i.e., IPCW), but the marginal survival probability for censoring at time $t$ is estimated using the CG estimator under an approximated Archimedean copula, parameterized by $\theta$. When $\theta\approx 0$, it yields the independence copula, which is equivalent to using the KM estimator as in Uno et al. [2011]. Our proposed concordance index thus requires a known copula, that models the dependency between marginals $T_{E}$ and $T_{C}$, but does not require interaction with the covariates. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, the traditional CI is computed as:
@@
Now, employing the IPCW technique used by Uno et al. [2011], but under dependent censoring, let $\hat{G}_{\text{CG}}(t)=\text{Pr}(T>t)$ represent the marginal survival probability for censoring at $t$ estimated by the CG estimator. We can then reweight Equation 3.1 using the CG estimator as:
@@
IBS-Dep: We propose a modified version of the Brier Score [Graf et al., 1999] under dependent censoring using margin time estimates as the censoring weights. The margin or “best-guess” time can be interpreted as a conditional expectation of the event time given the event time is greater than the censoring time. Given an instance is censored at time $t_{i}$, we can calculate its margin time using Equation 2.5, but with the CG estimator instead of the KM, i.e.$t_{e}^{\text{margin}}(t_{i})=t_{e}^{\text{margin}}(\hat{S}_{CG},t_{i})$, where $\hat{S}_{CG}$ is the marginal survival probability estimated using CG. We then formulate the dependent Brier Score as:
@@
$\text{where}\quad\tilde{e}_{i}=\begin{cases}t_{i}&\text{if}\quad\delta_{i}=1\\ t_{e}^{\text{margin}}(t_{i})&\text{elsewhere.}\end{cases}$
@@
MAE-Dep: We propose a modified version of the MAE-Margin [Haider et al., 2020] under dependent censoring. Again, we make use of the CG estimator to calculate margin times for the censored instances (Equation 2.5) and replace the censored event times with their “best-guess”-estimates. Furthermore, for each censored instance, we adopt the approach from Haider et al. [2020] and assign a confidence weight $\omega_{i}=1-S_{\text{\text{CG}}}(t_{c,i})$ to the error based on the margin value estimated using the CG method. For uncensored subjects, we set $w_{i}=1$. The dependent MAE is thus:
$\text{MAE}_{\text{Dep}}\left(\mathcal{D};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\Big{|}\big{[}(1-\delta_{i})\cdot t_{e}^{\text{margin (CG)}}(t_{i})$
@@ 4.1 Experimental Setup
Figure 2: Outline of the proposed evaluation pipeline. Based on some raw, survival dataset $\mathcal{D}$, we (1) form $\tilde{\mathcal{D}}$ by flipping the event bit of $\mathcal{D}$ and then estimate the feature-dependent censoring distribution, $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$. We then (2) select a subset of features from $\mathcal{D}$ using some strategy (e.g., Top-K features), and (3) generate a ground-truth dataset, $\mathcal{D}^{\prime}$, by removing the censored instances from $\mathcal{D}$. We then (4) create $\mathcal{D}^{\prime\prime}$ based on the ground-truth dataset, the learned censoring distribution, and the selected features, $\{X_{j}\}_{j\in\mathcal{S}}$. Using $\mathcal{D}^{\prime\prime}$, we (5) train two copula models, i.e., Clayton and Frank, and use AIC to select the copula $C_{\theta}$ that best suits the data. At the same time, we (6) train 5 survival learners independent of the copula on $\mathcal{D}^{\prime\prime}$ to produce survival predictions. Lastly, we calculate the (censored) evaluation error and (7) compare this to the true error.
@@
Synthetic datasets. Following the approach of Foomani et al. [2023], we create a synthetic survival dataset from a data generation process (DGP). The event and censoring distributions are specified by linear Weibull marginal distributions. Specifically, we sample $T_{E}$ and $T_{C}$ random variables from a known copula given Kendall’s $\tau$; if $\tau=0$, the marginals are independent, otherwise if $\tau>0$, the marginals are from an Archimedean copula, e.g., Clayton or Frank, known at evaluation time. See Appendix F.1 for more details.
@@
Semi-synthetic datasets. To evaluate our proposed metrics, we need to know the true metric, which requires precise knowledge of when each instance experiences the event. Due to censoring, event times are not available in real-world datasets for all instances, so we create a realistic, semi-synthetic dataset, where the covariates, event distribution, and censoring distribution align with some real-world datasets. This approach is inspired by Qi et al. [2023a], who proposed a semi-synthetic framework to generate different kinds of censoring distributions. However, in our case, we propose several strategies instead to create dependent censoring. In summary, our approach is as follows (see also the flowchart in Figure 2):
  1. 1.
    Start with some real-world survival dataset $\mathcal{D}$.
  2. 2.
    Based on some feature selection strategy (e.g., Top-$k$ features), select the relevant features from the original dataset $\mathcal{D}$ based on some technique, e.g., feature importances.
  3. 3.
    Calculate the censoring distribution of $\mathcal{D}$ to generate censor times, $\tilde{t}_{c,i}\sim G_{\text{CoxPH}(\tilde{\mathcal{D}})}(t\mid x_{i}$), where $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$ is the feature-dependent censoring distribution estimated by the Cox Proportional Hazard model [Cox, 1972] with Breslow estimator [Breslow, 1975].
  4. 4.
    Create $\mathcal{D}^{\prime}$ by removing the censored instances from $\mathcal{D}$, leaving just the uncensored ones.
  5. 5.
    Form $\mathcal{D}^{\prime\prime}$ by applying some reasonable, but synthetic censoring distribution to $\mathcal{D}^{\prime}$ and restrict the feature set to induce dependent censoring. Here we use the censoring distribution calculated in step 3 and the features calculated in step 2.
@@
Based on $\mathcal{D}$, we calculate the censoring distributions $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$ and apply it to $\mathcal{D^{\prime}}$, so that for each instance, that instance is synthetically censored if the censoring time is earlier than the event time ($\tilde{t}_{c,i}<t_{i}$), and otherwise is left uncensored. To induce dependent censoring, we consider three different feature selection strategies:
  • Original, where the original feature set is kept intact. This approach does not introduce additional residual dependence, but conditional independence may still not hold if important features are omitted originally.
@@
  • Top-$k$, where we select the $k$ most important features from $\mathcal{D}$ based on their predictive power. For this, we train a CoxPH model on $\mathcal{D}$ and then use permutation feature importances [Breiman, 2001] to select the $k$ best features. Feature importances are calculated by shuffling feature values and observing how much the model’s performance degrades, using Harrell’s CI.
@@
Copula model. To validate our dependent metrics in semi-synthetic datasets, we train two copula models, Clayton and Frank, and two nonlinear Weibull proportional-hazards models for the marginal distributions, $T_{E}$ and $T_{C}$, respectively, on $\mathcal{D}^{\prime\prime}$. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture with a single hidden layer and the nonlinear ReLU activation function. For each copula model, we train the marginals and the copula $C_{\theta}$ jointly using maximum likelihood estimation (MLE) (see Appendix F.3 for details). After training the two candidate models, we choose the model that minimizes the Akaike Information Criterion (AIC), given its loss and number of parameters. This offers a practical, data-driven approach to copula selection [Emura and Chen, 2018, Ch. 4]. We then use the best copula $C_{\theta}$ for the evaluation of the dependent metrics.
@@
Figure 3: Plot of the bias – i.e., the proximity to the true error – of established baseline metrics and the proposed dependent metrics as a function of the dependence under a known copula. We plot the mean biases across 10 experiments using different random seeds with data generated from a DGP ($N=10,000$, $d=10$). Overall, our dependent metrics show lower biases when dependency is present ($\tau>0.25$) than the independent metrics. See Appendix F.1 for details.
Given the semi-synthetic dataset $\mathcal{D}^{\prime\prime}$, we train the five survival learners across 5 experiments given a random seed. In each experiment, we follow Sechidis et al. [2011] and first split the dataset into training (70%), validation (10%), and test (20%) sets in a stratified manner wrt the event indicators and event times. This ensures a balanced representation of the time-to-event information and outcome across the data splits. After splitting, we impute missing values using sample mean for real-valued covariates or mode for categorical covariates based on the training set only. We then normalize numerical values and encode categorical features using a one-hot encoding strategy.
@@ 4.2 Results and Takeaways
Synthetic data. Figure 3 shows the evaluation bias – i.e., the proximity to true error – for four baseline metrics (Harrell’s CI, Uno’s CI, IPCW-integrated Brier Score, and MAE-Margin) and their dependent variants as a function of Kendall’s $\tau$. Assuming a known ground-truth copula (Archimedean family and dependency strength), we estimate $\hat{S}(t)$ using a CoxPH model to predict $N$ ISDs, where the risk scores are the negative predicted survival times. Overall, our dependent metrics show lower bias – i.e., closer proximity to the true error – when dependency is present ($\tau>0.25$). For the ranking metrics (Harrell’s CI, Uno’s CI), biases are similar under low dependency ($\tau<0.25$) but improve significantly under high dependency ($\tau>0.5$). For accuracy metrics (IPCW-integrated Brier Score, MAE-Margin), our dependent variants consistently provide better error estimates regardless of dependency or copula family. Discrepancies are most notable under a Clayton copula with moderate dependency ($0.50<\tau<$ 0.5).
As a practical example, consider a cancer study, with high dependent censoring ($\tau=0.75$) from a Clayton copula between event time $T_{E}$ (cancer relapse) and censoring time $T_{C}$ (cancer death). Based on Figure 3, if the true CI of a CoxPH model is 0.65, then, if we evaluate the model using Harrell’s CI, we anticipate the CI to be 0.685 ($\pm$ 0.035). If we use CI-Dep, we anticipate the CI to be 0.675 ($\pm$ 0.025).
Semi-synthetic data. Table 1 presents the evaluation biases in 3 selected datasets: METABRIC, MIMIC-IV and SEER. The complete results for all 6 datasets are in G.1. In the small-sized METABRIC dataset, Harrell’s gives superior estimates of the true concordance index, and there is a performance degradation in using the dependent variant. We see that CI-Dep degrades the performance compared to Uno’s CI, in this case, while IBS-Dep improves on the IPCW version, and MAE-Margin and MAE-Dep give similar results. In the medium-sized MIMIC-IV dataset, Harrell’s CI is again better, but our IBS-Dep and MAE-Dep improve on the IBS-IPCW and margin MAE, respectively, in all cases, however, the pseudo MAE is the best (i.e., it has the lowest error). In the large-sized SEER dataset, there are few differences in discriminative performances for the three metrics, but our IBS-Dep improves on IPCW, while our MAE-Dep improves on pseudo and margin. More often than not, Harrell’s and Uno’s CI (IPCW) give better estimates than our CI-Dep in these datasets, while our IBS-Dep and MAE-Dep give notable improvements to their independent counterparts. See Appendix G.1 for additional results.
@@
Table 1: Comparison of established survival metrics and the proposed dependent metrics in terms of the mean ($\pm$ SD.) bias, i.e., the proximity to the true error, averaged over 5 learners and 5 experiments. CI and IBS results are multiplied by 100.
In general, a (censored) metric should be representative of the true metric, i.e., it should tell us which model makes the best survival predictions – for example, by identifying the top-3 learners in alignment with the true ranking. Figure 4 shows the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the ordering of the top-3 survival learners by the true metric, over 5 experiments in the METABRIC dataset. We see general alignment between the metrics, but CI-Dep and IBS-Dep again show improvements over their independent variants.
Figure 4: Plot the ranking accuracy, i.e., the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the true ordering of the top-3 survival learners, out of 5 experiments in the METABRIC dataset.
@@ 5 Discussion
How Well Do The Metrics Work? As we do not have a formal proof of identifiability – i.e., we do not know the true relationship between $T_{E}$ and $T_{C}$ – we therefore study this question empirically by approximating $C_{\theta}$ and making a comparison between conventional survival metrics and our dependent variants. While the advantages to adopting our dependent metrics are evident in synthetic data, where we know the ground-truth copula and dependency level, these advantages vary with the dataset and its underlying distribution in the semi-synthetic datasets. Based on our observations, there is little to no improvement in adjusting the IPCW estimates using the CG estimator for discrimination (i.e., ranking) tasks. Concerning prediction tasks, our results show that it is better to use the CG estimator when (a) the dataset has a significant amount of censoring, approx. 20-25% or more, and (b) the dataset has an adequate size ($\geq$5,000 instances, like the MIMIC-IV and SEER cohorts).
How Do We Select the Copula? Estimation of the copula is inherently difficult due to the non-identifiability of competing risks data [Tsiatis, 1975], however, there exist several strategies to approximate the copula beside the likelihood-approach that we adopt: (a) cross-validation using Harrell’s CI [Emura and Chen, 2016], (b) simulation based on causal inference [Parikh et al., 2022, Mahajan et al., 2024], or (c) using domain knowledge to select the copula a priori [Foomani et al., 2023]. In (a), we first divide $N$ instances into $K$ groups of approximately equal sample sizes, and for each instance $i\in K$, we define a prognostic index, $PI_{i}(C_{\theta},\bm{x}_{i},W)$, which predicts the outcome for $i$, where $W$ are the model parameters. Using cross-validation, we pick the copula $\hat{\theta}$ which maximizes Harrell’s CI over possible values of Kendall’s $\tau$. However, this approach is computationally expensive. In (b), we generate counterfactual data with characteristics similar to the observed data and evaluate the model performance on the simulated data using counterfactual metrics (such as the Survival-$\ell_{1}$ [Foomani et al., 2023]). This approach is less computationally demanding, but requires adequate data and careful hyperparameter tuning [Parikh et al., 2022]. Finally, we can (c) plot the event and censoring distributions to visually inspect the curves. For example, when the censoring survival curve declines faster than the event curve, the event quantile exceeds the censoring quantile at a given time, indicating that a stronger dependence between the event and censoring times increases the likelihood of censoring before the event.

### v2

@@ 3 The Proposed Metrics
CI-Dep: We propose a modified version of Harrell’s CI [Harrell Jr et al., 1982] that employs inverse probability weighting (i.e., IPCW), but the marginal survival probability for censoring at time $t$ is estimated using the CG estimator under an approximated Archimedean copula, parameterized by $\theta$. When $\theta\approx 0$, it yields the independence copula, which is equivalent to using the KM estimator as in Uno et al. [2011]. Our proposed concordance index thus requires a known Archimedean copula, that models the dependency between marginals $E$ and $C$, but does not require interaction with the covariates. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, the traditional CI is computed as:
@@
Now, employing the IPCW technique used by Uno et al. [2011], but under dependent censoring, let $\hat{G}_{\text{CG}}(t)=\text{Pr}(C>t)$ represent the marginal survival probability for censoring at $t$ estimated by the CG estimator. We can then reweight Equation 3.1 using the CG estimator as:
@@
IBS-Dep: We propose a modified version of the Brier Score [Graf et al., 1999] under dependent censoring using margin time estimates as the censoring weights. The margin or “best-guess” time can be interpreted as a conditional expectation of the event time given the event time is greater than the censoring time. Given an instance is censored at time $t_{i}$, we can calculate its margin time using Equation 2.5, but with the CG estimator instead of the KM, i.e.$e^{\text{margin}}(t_{i})=e^{\text{margin}}(\hat{S}_{\text{CG}},t_{i})$, where $\hat{S}_{\text{CG}}$ is the marginal survival probability estimated using the CG estimator. We then formulate the dependent Brier Score as:
@@
$\text{where}\quad\tilde{e}_{i}=\begin{cases}e_{i}&\text{if}\quad\delta_{i}=1\\ e_{i}^{\text{margin}}(t_{i})&\text{elsewhere.}\end{cases}$
@@
MAE-Dep: We propose a modified version of the MAE-Margin [Haider et al., 2020] under dependent censoring. Again, we make use of the CG estimator to calculate margin times for the censored instances (Equation 2.5) and replace the censored event times with their “best-guess”-estimates. Furthermore, for each censored instance, we adopt the approach from Haider et al. [2020] and assign a confidence weight $\omega_{i}=1-\hat{S}_{\text{\text{CG}}}(t_{i})$ to the error based on the margin value estimated using the CG method. For uncensored subjects, we set $w_{i}=1$. The dependent MAE is thus:
$\text{MAE}_{\text{Dep}}\left(\mathcal{D};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\Big{|}\big{[}(1-\delta_{i})\cdot e^{\text{margin (CG)}}(t_{i})$
@@ 4.1 Experimental Setup
Figure 2: Outline of the proposed evaluation pipeline. Based on some raw, survival dataset $\mathcal{D}$, we (1) form $\tilde{\mathcal{D}}$ by flipping the event bit of $\mathcal{D}$ and then estimate the feature-dependent censoring distribution, $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$. We then (2) select a subset of features from $\mathcal{D}$ using some strategy (e.g., Top-K features), and (3) generate a ground-truth dataset, $\mathcal{D}^{\prime}$, by removing the censored instances from $\mathcal{D}$. We then (4) create $\mathcal{D}^{\prime\prime}$ based on the ground-truth dataset, the learned censoring distribution, and the selected features. Using $\mathcal{D}^{\prime\prime}$, we (5) train two copula models, i.e., Clayton and Frank, and use AIC to select the copula $C_{\theta}$ that best suits the data. At the same time, we (6) train 5 survival learners independent of the copula on $\mathcal{D}^{\prime\prime}$ to produce survival predictions. Lastly, we calculate the (censored) evaluation error and (7) compare this to the true error.
@@
Synthetic datasets. Following the approach of Foomani et al. [2023], we create a synthetic survival dataset from a data generation process (DGP). The event and censoring distributions are specified by linear Weibull marginal distributions. Specifically, we sample R.V.s $E$ and $C$ from a known copula given Kendall’s $\tau$; if $\tau=0$, the marginals are independent, otherwise if $\tau>0$, the marginals are from an Archimedean copula, e.g., Clayton or Frank, known at evaluation time. See Appendix F.1 for more details.
@@
Semi-synthetic datasets. To evaluate our proposed metrics, we need to know the true metric, which requires precise knowledge of when each instance experiences the event. Due to censoring, event times are not available in real-world datasets for all instances, so we create a realistic, semi-synthetic dataset, where the covariates, event distribution, and censoring distribution align with some real-world datasets. This approach is inspired by Qi et al. [2023a], who proposed a semi-synthetic framework to generate different kinds of censoring distributions. Figure 2 shows the pipeline to the create semi-synthetic datasets.
@@
In summary, based on some real-world survival dataset $\mathcal{D}$, we calculate the censoring distribution $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$, create $\mathcal{D^{\prime}}$ by removing the censored instances from $\mathcal{D}$, and then apply the censoring distribution to $\mathcal{D^{\prime}}$, so that for each instance, that instance is synthetically censored if the censoring time is earlier than the event time ($\tilde{c}_{i}<t_{i}$), and otherwise is left uncensored. We perform a feature selection step on $\mathcal{D}$ to induce dependent censoring using one of three strategies:
  • Original, where the original feature set is kept intact.
@@
  • Top-$k$, where we select the $k$ most important features from $\mathcal{D}$ based on their predictive power. We train a CoxPH model on $\mathcal{D}$ and then use permutation feature importances [Breiman, 2001] to select the $k$ best features. Feature importances are calculated by shuffling feature values and observing how much the model’s performance degrades, using Harrell’s CI.
@@
Finally, we have $\mathcal{D}^{\prime\prime}$: a dataset with a reasonable, but synthetic censoring distribution with a restricted feature set and where the ground-truth event times are known. This will allow us to compare the proposed and existing (censored) metrics with the ground-truth metrics.
Copula model. To validate our dependent metrics in semi-synthetic datasets, we train two copula models, Clayton and Frank, and two nonlinear Weibull proportional-hazards models for the marginal distributions, $E$ and $C$, respectively, on $\mathcal{D}^{\prime\prime}$. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture with a single hidden layer and the nonlinear ReLU activation function. For each copula model, we train the marginals and the copula $C_{\theta}$ jointly using maximum likelihood estimation (MLE) (see Appendix F.3 for details). After training the two candidate models, we choose the model that minimizes the Akaike Information Criterion (AIC), given its loss on the validation set and number of parameters. This offers a practical, data-driven approach to copula selection [Emura and Chen, 2018, Ch. 4] by selecting the copula model with the lowest validation loss. We then use the best copula $C_{\theta}$ for the evaluation of the dependent metrics.
@@
We train the five survival learners across 5 experiments given a random seed (0-4). In each experiment, we follow Sechidis et al. [2011] and first split the dataset into stratified training (70%), validation (10%), and test (20%) sets. After splitting, we impute missing values using sample mean for real-valued covariates or mode for categorical covariates based on the training set only. We then normalize numerical values and encode categorical features using a one-hot encoding strategy. All results are reported in the test sets.
Figure 3: Plot of the bias ($Y$-axis) – i.e., the proximity to the true error – of established baseline metrics and the proposed dependent metrics as a function of the dependency ($X$-axis) under a known copula. We plot the mean biases ($\pm$ SD.) in the test sets across 100 experiments in a synthetic dataset ($N=10,000$, $d=10$). Overall, our dependent metrics show lower biases when dependency is notable ($\tau>0.25$) than the independent metrics. See Appendix F.1 for details.
@@ 4.2 Results and Takeaways
Synthetic data. Figure 3 shows the evaluation bias – i.e., the proximity to true errorc – for four baseline metrics (Harrell’s CI, Uno’s CI, IPCW-integrated Brier Score, and MAE-Margin) and their dependent variants as a function of Kendall’s $\tau$. Assuming a known ground-truth copula (Archimedean family and dependency strength), we estimate $\hat{S}(t)$ using a CoxPH model to predict $N$ ISDs, where the risk scores are the negative predicted survival times. Overall, our dependent metrics show lower bias – i.e., closer proximity to the true error – when dependency is notable ($\tau>0.25$). For the ranking metrics (Harrell’s CI, Uno’s CI), biases are similar under low dependency ($\tau<0.25$) but improve significantly under high dependency ($\tau>0.5$). For accuracy metrics (IPCW-integrated Brier Score, MAE-Margin), our dependent variants consistently provide better error estimates regardless of dependency or copula family.
As a practical example, consider a cancer study, with high dependent censoring ($\tau=0.75$) from a Clayton copula between event time $E$ (cancer relapse) and censoring time ${C}$ (cancer death). Based on Figure 3, if the true CI of a CoxPH model is 0.65, then, if we evaluate the model using Harrell’s CI, we anticipate the CI to be 0.68 ($\pm$ 0.03). If we use CI-Dep, we anticipate the CI to be 0.67 ($\pm$ 0.02).
Semi-synthetic data. Table 1 presents the results in 3 semi-synthetic datasets: METABRIC, MIMIC-IV and SEER. The complete results for all 6 datasets are in Appendix G.1. In the small-sized METABRIC dataset, Harrell’s gives superior estimates of the true concordance index, and there is a performance degradation in using the dependent variant. We see that CI-Dep degrades the performance compared to Uno’s CI, in this case, while IBS-Dep improves on the IPCW version, and MAE-Margin and MAE-Dep give similar results. In the medium-sized MIMIC-IV dataset, Harrell’s CI is again better, but our IBS-Dep and MAE-Dep variants improve on the IBS-IPCW and margin MAE, respectively, in all cases, however, the pseudo MAE is the best (i.e., it has the lowest error). In the large-sized SEER dataset, there are few differences in discriminative performances for the three metrics, but our IBS-Dep improves on IPCW, while our MAE-Dep improves on pseudo and margin. More often than not, Harrell’s and Uno’s CI (IPCW) give better estimates than our CI-Dep in these datasets, while our IBS-Dep and MAE-Dep give notable improvements to their independent counterparts. See Appendix G.1 for additional results.
@@
Table 1: Comparison of established survival metrics and the proposed dependent metrics in terms of the mean ($\pm$ SD.) bias in the semi-synthetic test sets, averaged over 5 learners and 5 experiments. CI and IBS results are multiplied by 100.
In general, a (censored) metric should be representative of the true metric, i.e., it should tell us which model makes the best survival predictions. Therefore, we create a new experiment as follows:
  • 1.
    In our semi-synthetic datasets, we are aware of the ground-truth (the actual event times) of all instances. For each model, we use this information to calculate the True CI, True IBS and True MAE results – that is, how good are the survival predictions using each metric, if all the instances were observed (not censored).
  • 2.
    For each true metric, we then rank the top-3 survival models – e.g., CoxPH, RSF and DeepSurv – by their true metric. In this example, CoxPH has the highest True CI, then RSF and then DeepSurv.
  • 3.
    For each censored metric (proposed and existing), we check if the metric gives the same ordering of survival models as its corresponding true metric – e.g., if True CI indicates that CoxPH, RSF and DeepSurv are the three best models, does CI-Dep indicate the same? This experiment is repeated 5 times with different data splits.
Figure 4 shows the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the ordering of the top-3 survival learners in the METABRIC test set. See Appendix G for results in all semi-synthetic datasets.
Figure 4: Plot the ranking accuracy, i.e., the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the true ordering of the top-3 survival learners, out of 5 experiments in the METABRIC test sets.
@@ 5 Discussion
How do the Metrics Work? As we do not have a formal proof of identifiability – i.e., we do not know the true relationship between $E$ and $C$ – we therefore study this question empirically by approximating $C_{\theta}$ and making a comparison between conventional metrics and our dependent variants. While the advantages to adopting our dependent metrics are evident in synthetic data, where we know the ground-truth copula and dependency level, these advantages vary with the dataset and its underlying distribution in the semi-synthetic datasets. Based on our observations, the most consistent trend is seen in the proposed MAE-Dep –- which is directly comparable to the MAE-Margin. Specifically, the MAE-Dep always outperforms its independent counterpart when censoring exceeds 20%. In these scenarios, MAE-Dep yields lower mean biases across five repeated experiments on all four dataset variants with over 20% censoring. This improvement is encouraging, as it holds across datasets with possibly varying characteristics and censoring mechanisms. However, when the censoring rate falls between 0% and 20%, the trend reverses slightly: the MAE-Dep tends to be marginally higher than the MAE-Margin.
What are the Limitations? Our metrics rely on a known or approximate copula to adjust for dependent censoring, but estimating the copula is inherently difficult due to the non-identifiability of competing-risks data [Tsiatis, 1975]. This poses a key limitation for applying our method to real-world datasets. However, there exist several strategies to approximate the copula using observed data: Emura and Chen [2016] proposed to divide the $N$ instances into $K$ groups of approximately equal sample sizes, and for each instance $i\in K$, define a prognostic index, $\text{PI}_{i}(C_{\theta},\bm{x}_{i},W)$, which predicts the outcome for $i$, where $W$ are the model parameters. Using cross-validation, we pick the copula $\hat{\theta}$ which maximizes the CI over possible values of Kendall’s $\tau$. Parikh et al. [2022] and Mahajan et al. [2024] proposed generating counterfactual data with characteristics similar to the observed data and evaluate the model performance on the simulated data using counterfactual metrics (such as the Survival-$\ell_{1}$ by Foomani et al. [2023]). Foomani et al. [2023] proposed to first get an intuition about the underlying copula (e.g., Clayton, Frank) by plotting the event and censoring distributions to visually inspect the curves and then go on to approximate $C_{\theta}$ using MLE.